UT Sep 04, 2019 - Restart Night #3


OSA: S. Allanson

ISp: E. Solheid

Mtn Manager: P. Hartley

ISA: B. Rothberg (Tucson), O. Kuhn (Tucson)

AO: D. Miller (Tucson)

Observers/Technical Expert: J. Hill (Tucson)



Rollover of previous work
  • MODS On-Sky Checkout
    • DONE Rotator Centers
      • Collimate and then send track presets to the same field at PA=0, 180 and PA=90, 270. Analyze data with modsrc . modsrc works on both MODS and LBC images and it computes rotator centers for pairs of images which do not have to be taken at diametrically opposed position angles but can be, say, 90 deg apart instead.
    • DONE Rotator Zero Points
      • Acquire 2 MOS masks and note the delta theta offset needed to align the mask. Ideally, the rotator zero point will be set (in PCSInstrument.conf) so that the mask or longslit is oriented at PA=0.
      • Finished Stone A for MODS-2 - theta=-0.018, M1 theta=-0.116, M2 theta (2nd time) = -0.023
      • Finished Stone O for both MODS, M1 theta=-0.145, M2 theta=-0.016
    • DONE Dither Test using Pointing Mask
    • DONE Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode
    • DONE Photometric Field Observations
    • DONE Test Science OB
      • DONE Sidereal
      • DONE Non-Sidereal

UT 02:16 - closed due to clouds and rain, running MODS-Bino cals (revised versions) - MODS-1 AGW WFS fixed (badly seated fiber)

UT 02:46 - heard thunder up there, still raining (and obviously still closed)

-IT 7285 occurred -mods1data froze mid-exposure (during cals)

UT 04:46 - readout artifact on MODS-1R #0022 (black rectangle) - appears to be a sequencer issue -will keep an eye out if it happens again (and still closed for rain/clouds)

UT 05:41 - might be improving

UT 06:07 - opening enclosure - will proceed to Stone O first to measure rotator ZP with astrometric mask

UT 06:22 - pointing and collimating near Stone O

UT 06:24 - sending preset to Stone O for rotator ZP
modsAlign mods1r.20190904.0035.fits mods1r.20190904.0036.fits mods.36.StoneO.mms
Computed Mask Alignment Offset:
   dX = -0.686 arcsec
   dY = 1.063 arcsec
  dPA = -0.145 deg

MODS1 Offset Command:
   offsetpointing -0.1452 -0.686  1.063 detxy rel

and

lbto@obs2:2 % modsAlign mods2r.20190904.0041.fits mods2r.20190904.0042.fits mods.36.StoneO.mms

Computed Mask Alignment Offset:
   dX = 3.792 arcsec
   dY = -1.471 arcsec
  dPA = -0.016 deg

MODS2 Offset Command:
   offsetpointing -0.0157  3.792 -1.471 detxy rel

UT 06:38 - proceeding to M52 to measure the rotator center

-Active preset, then track presets at pairs of 0,180 and 90,270 using pointing mask

UT 06:48 - sending PA=0 track preset UT 06:53 - sending PA=180 track preset UT 06:57 - sending PA=90 track preset UT 07:00 - sending PA=270 track preset

UT 07:07 - sending preset to Stone A for rotator ZP

lbto@obs2:32 % modsAlign mods1r.20190904.0043.fits mods1r.20190904.0044.fits mods.37.StoneA.mms

Computed Mask Alignment Offset:
   dX = 0.011 arcsec
   dY = 0.693 arcsec
  dPA = -0.116 deg

MODS1 Offset Command:
   offsetpointing -0.1159  0.011  0.693 detxy rel

lbto@obs2:77 % modsAlign mods2r.20190904.0049.fits mods2r.20190904.0050.fits mods.37.StoneA.mms

Computed Mask Alignment Offset:
   dX = 4.390 arcsec
   dY = -2.155 arcsec
  dPA = -0.023 deg

MODS2 Offset Command:
   offsetpointing -0.0233  4.390 -2.155 detxy rel

UT 07:24 - moving to spectrophotometric standard Feige 110

-had to nudge each acquisition -0.5 M1, -0.25 M2

UT 07:33 -starting spectroscopic observations - going through all modes

UT 08:28 - completed Feige 110

UT 08:31 - NS sidereal test Bino-MODS imaging 2018 YH (C. Veillet)

-can see it in MODS2, but it landed on a bad column on MODS2R

UT 08:49 - sending an async offset -10" +15" for MODS2

-both offsets show as absolute det x y

The offset was sent through the MODS GUI, and the dropdown menus for CoordSys and MoveType were already showing "DETXY" and "Relative" so these were not touched. In the ISIS log, however, the offset does show up as an absolute offset:

2019-09-04T08:49:30.776295 MC2>M2.TC offsetpointing 0 -10 15 DETXY ABS right

I am adding a note after the fact because I think I understand why the offset I made through the GUI was ABS and not REL. The script was sending absolute offsets and I tried to nudge the pointing while it was running. The GUI does not update dynamically (will have to test whether clicking "update" will cause the offset CoordSys and MoveType to change) and it was probably misrepresenting the MoveType. The moral of the story: if you really want to make sure you are sending the type of offsets you think you are sending, select the desired CoordSys & MoveType from the GUI.

UT 09:04 - long lag on MODS-1 Blue

UT 09:05 - completed 2018 YH

UT 09:06 - moving to spec phot std BD+28 - all modes

UT 09:23 - another artifact mods1r.20190904.0069 - central rectangle around spectra looks like a sequencer problem

UT 09:51 - need to do a blue fitsflush MODS-1

UT 10:11 - Moving to a test Science OB - NGC 1110

MODS2 Computed Slit Alignment Offset:
dX = 5.137 arcsec
dY = 10.777 arcsec

MODS1 Computed Slit Alignment Offset:
dX = 0.692 arcsec
dY = 14.430 arcsec
-aligned on brightest blob/star in acquisition images:
MODS1: Object centroid: X=1522.27+/-0.08 Y=1531.53+/-0.11
MODS2: Object centroid: X=1494.47+/-0.05 Y=1534.24+/-0.09
UT 10:28 - starting spectroscopic observations
DIMM 0.72", WFS Guiders 0.7-0.6"
Airmass 1.33

Taking a 2nd set of exposures 600sec (total 1200sec each MODS) DIMM 0.66"

UT 10:54 - done

UT 10:55 - moving to SDSS photometric field

-target too low - Shells RIP'd

UT 11:00 - recovered from RIP

UT 11:07 - will move to M52 dither test

-was a track script - correcting

-moving to a standard field instead RA 02:33:37 Dec -05:17:48 (PG0231)

-reducing exposure time

UT 11:52 - finished specstd - doing a dithering test now nearby

UT 12:02 - MODS-2 exposed while moving M2R 0102, M2B 0088 - race condition

UT 12:06 - MODS-2 fixed, taking data again

UT 12:16 - MODS1R is lagging behind the other channels

UT 12:18 - stopping script

Starting skyflats

-accidentally set to 1kx1k

-too bright for MODS-2 will attempt a few more for MODS-1

UT 12:44 - putting MODS to sleep

-- BarryRothberg - 03 Sep 2019
Topic revision: r11 - 05 Sep 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback